# How does Transformer Learn Implicit Reasoning?

### v1

@@ 1 Introduction
Large language models (LLMs) demonstrate strong performance on complex, multi-step reasoning tasks [9, 7, 27, 1, 31, 18]. Typically, these reasoning abilities are elicited using chain-of-thought (CoT) prompting, which encourages models to explicitly articulate intermediate reasoning steps [26, 34, 6, 32, 28]. Beyond CoT, recent studies indicate that LLMs can also engage in implicit reasoning [36, 5, 11, 14], producing correct answers without verbalizing intermediate steps.
@@
Existing studies that investigate implicit reasoning often rely on pretrained LLMs whose training data lacks precise experimental control, making it challenging to conclusively determine whether models have genuinely learned implicit multi-step reasoning or instead rely on its prior knowledge or shortcut solutions [12, 33, 5]. Symbolic datasets [24, 25] partially alleviate this concern by training models from scratch, yet they still lack the fine-grained experimental control and behavioral granularity necessary for deeper analysis. To address these limitations, we construct an extended symbolic environment, featuring targeted omissions and query-level variations, to precisely identify whether implicit reasoning and generalization truly emerge.
@@
Figure 1: We track the evolution of bridge entity representations across training. In the memory phase (left), intermediate entities are explicitly decodable but geometrically scattered. In the implicit reasoning phase (right), representations converge in cosine space, supporting structured multi-hop reasoning through early-to-late layer transitions. (a) Illustration of how shared-immediate-entity queries are composed to form 2-hop queries used in cosine-based representation analysis.
@@
In Section 4, we close the loop by explicitly connecting these internal representational mechanisms to external behavioral patterns. We demonstrate that successful generalization robustly correlates with the clustering structure of intermediate representations across diverse queries and training distributions. Although ID triple supervision is not required to induce this clustering, it substantially accelerates its emergence by constraining the representational space early in training. Finally, we identify that what appears to be first-hop generalization to out-of-distribution (OOD) triples is actually an artifact arising from representational alignments induced by ID exposure, highlighting the fragile and data-dependent nature of implicit generalization.
@@ 2 Behavioral Signatures of Implicit Reasoning under Fine-Grained Control
Existing studies on implicit reasoning fall into two broad categories, each with notable limitations. (1) Analyses based on pretrained LLMs operate in an uncontrolled setting where the training data are opaque—making it difficult to distinguish genuine reasoning from memorization. (2) In contrast, recent works adopt symbolic datasets with synthetic training from scratch [24], but primarily focus on dataset-level trends, without isolating what specific training signals are necessary for solving each compositional query.
@@ 2.1 Data Construction: Fine-Grained Control for Compositional Reasoning
To enable fine-grained behavioral analysis, we extend the symbolic reasoning setup of Wang et al. [24] with expressive query-level control configurations. The data comprises atomic triples and compositional queries:
  • Atomic Triples. Each atomic fact is represented as a triple $(e_{1},r_{1})\rightarrow e_{2}$, partitioned into two subsets: In-Distribution (ID) Triples are used in both standalone form and as components of multi-hop training queries; Out-of-Distribution (OOD) Triples appear in training data but are excluded from multi-hop construction, enabling the creation of test queries involving out-of-distribution reasoning. Note that ID and OOD triples share the same set of entities and relations.
  • 2-Hop Queries. Each reasoning task takes the form of a compositional chain $(e_{1},r_{1},r_{2})\rightarrow e_{3}$, where the model performs implicit reasoning over an intermediate entity. We distinguish: Train-II: queries with both hops from ID triples used during training; Test-OI: test queries where the first hop comes from an OOD triple and the second hop from an ID triple. Other query types, such as Test-II, Test-IO, and Test-OO, follow similar definitions.
@@ 2.2 Three-Stage Generalization
Phase II: ID Generalization. After memorization saturates, the model begins to generalize to Test-II queries (unseen ID-ID compositions), marking a shift from memorization to compositional generalization within ID, akin to the grokking phenomenon described by Wang et al. [24].
@@ 2.3 ID Triples Are Not Required for ID Generalization—but Accelerate It
Prior work has repeatedly observed that while models can correctly answer individual atomic triples, they often fail to generalize to 2-hop queries constructed by composing those same triples[3, 29, 41]. Additionally, in Section 2.2, we observe that our model quickly memorized atomic triples (Phase I) but took longer to generalize to Test-II queries. These findings raise a natural question: Are atomic ID Triples actually necessary for learning ID-based 2-hop reasoning?
@@ 3 Locating and Characterizing Reasoning Representations
In Section 2, we observed surprising behavioral phenomena that offer new insights into implicit reasoning in Transformers. These findings challenge prevailing assumptions, such as the necessity of exposure to atomic facts [39, 35] or the impossibility of OOD generalization [24].
@@
3.1 Locating Intermediate Entity Representations via Cross-Query Semantic Patching
To analyze how transformers internally represent intermediate reasoning steps, a crucial first step is to identify where such representations are encoded in the model’s hidden states. Existing methods such as linear probing and causal patching offer only partial insight. Linear probing reveals correlations between hidden states and output tokens, but not their causal role in reasoning. Causal patching assesses causal influence, typically measures whether a random source activation affects the target’s output, but doesn’t assess what the activation semantically represents.
Cross-Query Semantic Patching. To provide deeper causal interpretation, we propose cross-query semantic patching, which examines the specific semantic drift after causal patching. Specifically, given a source query $(e_{1},r_{1},r_{2})$, we extract the hidden representation believed to encode the bridge entity from a selected position(i.e., layer 3 at $r_{1}$ position), and then insert it into a structurally similar target query $(e_{5},r_{6},r_{7})$ at the same position. This hidden state is hypothesized to represent the intermediate entity $r_{1}(e_{1})$, which in the source query leads to the final answer $r_{2}(r_{1}(e_{1}))$. In contrast, the unpatched target query normally follows the path $r_{7}(r_{6}(e_{5}))$. If patching results in $r_{7}(r_{1}(e_{1}))$, we conclude that the hidden state encodes a semantically valid intermediate entity, suggesting it is both causally relevant and semantically reusable across queries.
@@ 3.1 Locating Intermediate Entity Representations via Cross-Query Semantic Patching
To systematically identify where such representations reside, we first construct a wide range of query pairs, and then apply the patching procedure across multiple layers and token positions. A patching attempt is considered successful only if the model’s output exactly matches the expected answer implied by the inserted intermediate entity, and we compute the success rate for each candidate location accordingly.
We perform this analysis under three settings that differ in both the source of the intermediate entity and the model’s training stage: (1) Phase II with ID-derived intermediate entities, (2) Phase III with ID-derived intermediate entities, and (3) Phase III with OOD-derived intermediate entities. This alignment ensures that patching is conducted under conditions where the model is capable of reasoning over the relevant intermediate entity type. For completeness, we also report Phase I results, where patching yields negligible success, confirming that reasoning-relevant representations emerge only after generalization.
@@ 3.2 Explicitly Decodable $\neq$ Implicitly Informative
Having located the positions encoding intermediate entities, we next ask whether their internal role can be explained using existing interpretability tools. A key assumption in prior work is that if a hidden state encodes an intermediate entity, it should be decodable into a human-interpretable token, for example via the logit lens [17, 15, 32, 21, 39]. We test this assumption by measuring the decodability of $\mathbf{h}_{\texttt{r}_{1}}^{5}$, and examining whether decodability aligns with the emergence of reasoning behavior across training phases.
@@ 3.3 Geometric Regularity of Intermediate Representations via Cosine Lens
Most prior work focuses on decoding representations, but we take the reverse approach: given a known intermediate entity, can we identify recurring structure in how the model represents it across contexts to achieve consistent representations [23]? This reverse mapping is enabled by our earlier analysis in Section 3.1, which identifies the position $\mathbf{h}_{\texttt{r}_{1}}^{5}$ encoding the intermediate entity. At this anchor, we collect hidden states from queries sharing the same intermediate entity (Figure 1(a)), and examine whether these vectors reflect a consistent internal pattern.
To assess consistency, we focus on structural alignment in the model’s embedding space, using cosine similarity—a common metric for semantic proximity in high-dimensional representations[8, 20, 13]. This allows us to examine whether the model reuses internal abstractions through representational geometry, instead of relying on explicit decodability.
[…]
@@
Tracking these scores, we find that the ID Cohesion Score rises steadily, aligning with Test-II generalization, while the OOD Alignment Score begins to increase later, tracking the rise in Test-OI accuracy (Figure 2). This pattern indicates that successful implicit reasoning is closely associated with representational consistency across contexts: hidden states corresponding to the same intermediate entity become more geometrically aligned as training progresses. These clustering dynamics appear to play an important role in enabling the reuse of latent representations during multi-hop inference, suggesting that cosine-space structure may be a key internal signal supporting generalization.
@@ 4 Closing the Loop: Explaining Behavioral Phenomena
Drawing on the mechanistic evidence in Section 3, we revisit our empirical observations in Section 2 and explain how they emerges from the internal dynamics of representation formation.
4.1 Representation Clustering Is Necessary but Not Sufficient for Generalization
In Section 3.3, we observed a striking alignment between the emergence of representational clustering and behavioral generalization: as internal representations of intermediate entities became more geometrically coherent, accuracy on corresponding multi-hop queries improved. While this alignment is compelling, it does not by itself establish a complete explanatory link.
Such correspondence invites caution. Multi-hop reasoning requires two distinct capabilities: (1) forming a consistent internal representation of the intermediate entity, and (2) correctly using that representation to predict the final answer. Clustering supports the first step, but successful generalization depends on the second step as well. It is therefore possible that clustering is a necessary condition for generalization, but not by itself sufficient.
We hypothesize that in our base setup, the model learns representation formation and usage in parallel, causing ID Cohesion and Test-II accuracy to rise together—creating the illusion of a direct causal link. By contrast, in settings with higher model capacity or task complexity (e.g., larger models or 3-hop queries), this process decouples. As shown in Appendix G.2 and Appendix C, the ID Cohesion score stabilizes early, while ID generalization emerges only gradually. This indicates that forming consistent representations is not sufficient: the second-hop mapping must also be learned.
Phase III offers a clean testbed to validate this interpretation. In Test-OI queries, the second-hop relation is drawn from the ID set and is already well learned during Phase II. Consequently, generalization depends primarily on the success of the first step: whether the model can align OOD-derived intermediate representations with the ID-based cluster structure. As predicted, we observe a strong coupling between the OOD Alignment metric and Test-OI (Test-OII in the 3-hop setting) accuracy. This pattern confirms that once the second-hop mapping is in place, clustering alone enables generalization—supporting its role as a necessary but not sufficient condition in the reasoning pipeline.
@@ 4.4 Why the First Hop, and Only the First Hop, Generalizes to OOD?
[…]
@@ 5 Discussion
Our study, conducted in a controlled symbolic dataset environment, reveals key insights into the mechanisms of implicit reasoning in transformers, highlighting specific patterns and behaviors that clarify how multi-hop implicit reasoning emerges. These findings may provide valuable answers to existing questions about the implicit reasoning capabilities of LLMs. For instance, our observation regarding the requirement for query-level match offers a potential explanation for why knowledge learned from single-hop tasks does not easily transfer to multi-hop reasoning in LLMs [3, 38, 29]. However, it is important to note that LLMs operate with far richer and more complex knowledge bases, and their internal knowledge interaction mechanisms likely differ from those in our controlled environment. Therefore, while our findings offer useful insights, they should be regarded as preliminary guidance rather than a complete explanation of the reasoning dynamics in LLMs.
@@ Appendix A Related Work
The initial focus on the mechanisms of implicit reasoning arose from the discovery that traditional single-hop knowledge editing methods are ineffective in the context of multi-hop implicit reasoning [40, 37, 12]. This phenomenon since gained more significant attention, prompting researchers to investigate the mechanisms underlying implicit reasoning. Some works suggest that the failure of implicit reasoning is due to the intermediate entities not being properly processed [21, 32]; Other works suggest that the reason lies in the intermediate entities being processed, but not being passed to the correct position [15, 35, 5]. The parallel exploration paths within the model are also a research direction [22, 25]. Recent works have found that the different knowledge composed as different hops used in implicit reasoning is stored in different layers of the model [39, 35]. Additionally, some studies propose that models rely on shortcuts to successfully complete implicit reasoning [33].
@@
Some works have found that the conditions for models to learn implicit reasoning are quite demanding. Some have discovered through custom symbolic datasets that models generalize implicit reasoning abilities only after grokking, and that they require highly compositional data  [24]. Other works have found that simply providing optical knowledge is insufficient for enabling models to perform implicit reasoning, it requires training with corresponding multi-hop reasoning data [29, 38, 41]. Moreover, it has been observed that when the two pieces of foundational knowledge used for 2-hop implicit reasoning appear in different paragraphs of the training corpus, the model struggles to combine them for implicit reasoning. However, when these pieces appear within the same paragraph, the model’s accuracy in answering the corresponding 2-hop queries significantly improves [3].
@@
As for the probing tools, most of the work uses decoding-based methods to probe intermediate entities. Many works assume that if a model encodes an intermediate entity, it should be able to explicitly decode it. They use the logit lens [17]. as evidence of the presence of intermediate entities in implicit reasoning [21, 32, 15]. A few studies have also trained a linear transformation layer to decode intermediate entities [22].
@@
Although current Chain-of-Thought (CoT) based reasoning models have shown impressive performance, some perspectives argue that the thought chains do not truly reflect the model’s reasoning process [16, 2]. Some works have begun to explore Chain of Thought (CoT) methods that do not verbalize intermediate steps. One approach is to replace the Chain of Thought (CoT) with dots, adjusting the number of dots based on the original length of the thought chain [19, 4]. Another approach to avoid verbalizing is to bypass decoding the tokens in the CoT process and directly use the last hidden state from the previous step as the input vector for the next step [10]. These approaches have achieved results comparable to traditional CoT on specific tasks.

### v2

# How do Transformers Learn Implicit Reasoning?

@@ 1 Introduction
Large language models (LLMs) demonstrate strong performance on complex, multi-step reasoning tasks [deepseek-r1, gemini25pro, xai2025grok3, anthropic2025claude, qwen3, openai2025o3, DBLP:conf/icml/KhonaOHRNDLT24]. Typically, these reasoning abilities are elicited using chain-of-thought (CoT) prompting, which encourages models to explicitly articulate intermediate reasoning steps [cot-prompt, tot, lrm-survey, yang2024large1, atomr]. Beyond CoT, recent studies indicate that LLMs can also engage in implicit reasoning [system2-system1, hopping-too-late, think-prune, k1.5], producing correct answers without verbalizing intermediate steps.
@@
Existing studies that investigate implicit reasoning often rely on pretrained LLMs whose training data lacks precise experimental control, making it challenging to conclusively determine whether models have genuinely learned implicit multi-step reasoning or instead rely on its prior knowledge or shortcut solutions [ju2024investigating, yang2024large2, hopping-too-late]. Symbolic datasets [wang2024grokking, wang2024understanding, DBLP:conf/nips/0002SFST024] partially alleviate this concern by training models from scratch, yet they still lack the fine-grained experimental control and behavioral granularity necessary for deeper analysis. To address these limitations, we construct an extended symbolic environment, featuring targeted omissions and query-level variations, to precisely identify whether implicit reasoning and generalization truly emerge.
@@
Figure 1: We track the evolution of bridge entity representations across training. In the memory phase (left), intermediate entities are explicitly decodable but geometrically scattered. In the implicit reasoning phase (right), representations converge in cosine space, supporting structured multi-hop reasoning through early-to-late layer transitions. Queries that share an bridge entity are composed for cosine-based representation analysis.
@@
In Section 4, we close the loop by explicitly connecting these internal representational mechanisms to external behavioral patterns. We demonstrate that successful generalization robustly correlates with the clustering structure of intermediate representations across diverse queries and training distributions. Although in-distribution (ID) triple supervision is not required to induce this clustering, it substantially accelerates its emergence by constraining the representational space early in training. Finally, we identify that what appears to be first-hop generalization to out-of-distribution (OOD) triples is actually an artifact arising from representational alignments induced by ID exposure, highlighting the fragile and data-dependent nature of implicit generalization.
@@ 2 Behavioral Signatures of Implicit Reasoning under Fine-Grained Control
Existing studies on implicit reasoning fall into two broad categories, each with notable limitations. (1) Analyses based on pretrained LLMs operate in an uncontrolled setting where the training data are opaque—making it difficult to distinguish genuine reasoning from memorization. (2) In contrast, recent works adopt symbolic datasets with synthetic training from scratch [wang2024grokking], but primarily focus on dataset-level trends, without isolating what specific training signals are necessary for solving each compositional query.
@@ 2.1 Data Construction: Fine-Grained Control for Compositional Reasoning
To enable fine-grained behavioral analysis, we extend the symbolic reasoning setup of wang2024grokking with expressive query-level control configurations. The data comprises atomic triples and compositional queries:
  • Atomic Triples. Each atomic fact is represented as a triple $(e_{1},r_{1})\rightarrow e_{2}$. This formulation mimics simple factual relations such as (Alice, mother-of) → Beth and (Beth, sister-of) → Carol, serving as the atomic unit of the reasoning environment. The triples are partitioned into two subsets: In-Distribution (ID) Triples are used in both standalone form and as components of multi-hop training queries; Out-of-Distribution (OOD) Triples appear in training data only in standalone form, and are excluded from multi-hop composition, enabling the creation of test queries involving out-of-distribution reasoning. Note that ID and OOD triples share the same set of entities and relations.
  • 2-Hop Queries. Each reasoning task takes the form of a compositional chain $(e_{1},r_{1},r_{2})\rightarrow e_{3}$, where the model performs implicit reasoning over an bridge entity $e_{2}$. For instance, the model receives only the compositional query (Alice, mother-of, sister-of) and is expected to predict the correct target Carol, implicitly reasoning through the intermediate entity Beth. We distinguish: Test-OI: test queries where the first hop comes from an OOD triple and the second hop from an ID triple; Train-II: queries with both hops from ID triples used during training. Other query types, such as Test-II, Test-IO, and Test-OO, follow similar definitions.
@@ 2.2 Three-Stage Generalization
Phase II: ID Generalization. After memorization saturates, the model begins to generalize to Test-II queries (unseen ID-ID compositions), marking a shift from memorization to compositional generalization within ID, akin to the grokking phenomenon described by wang2024grokking.
@@ 2.3 ID Triples Are Not Required for ID Generalization—but Accelerate It
Prior work has repeatedly observed that while models can correctly answer individual atomic triples, they often fail to generalize to 2-hop queries constructed by composing those same triples[balesni2024two, xu2025memorizing, zhu2024towards]. Additionally, in Section 2.2, we observe that our model quickly memorized atomic triples (Phase I) but took longer to generalize to Test-II queries. These findings raise a natural question: Are atomic ID Triples actually necessary for learning ID-based 2-hop reasoning?
@@ 3 Locating and Characterizing Reasoning Representations
In Section 2, we observed surprising behavioral phenomena that offer new insights into implicit reasoning in Transformers. These findings challenge prevailing assumptions, such as the necessity of exposure to atomic facts [zhang2024locate, yao2025cake] or the impossibility of OOD generalization [wang2024grokking].
@@
3.1 Locating Intermediate Entity Representations via Cross-Query Patching
To analyze how transformers internally represent intermediate reasoning steps, a crucial first step is to identify where such representations are encoded in the model’s hidden states.
Existing methods such as linear probing and causal patching offer only partial insight. Linear probing reveals correlations between hidden states and output tokens, but not their causal role in reasoning. Causal patching assesses causal influence, typically measures whether a random source activation affects the target’s output, but doesn’t assess what the activation semantically represents [wang2024grokking].
Cross-Query Semantic Patching. To go beyond correlation or superficial causal influence, we introduce cross-query semantic patching, a method designed to test whether a hidden representation encodes a semantically valid intermediate entity. Specifically, given a source query $(e_{1},r_{1},r_{2})$, we test a set of candidate hidden states from different layers and positions (e.g., layer 3 at the $r_{1}$ position) that may contain the bridge entity representation. For each candidate, we insert its hidden vector into a structurally similar target query $(e_{5},r_{6},r_{7})$ at the same position, replacing the original hidden state.
@@ 3.1 Locating Intermediate Entity Representations via Cross-Query Patching
If the patched model’s prediction changes from the original reasoning path $r_{7}(r_{6}(e_{5}))$ to $r_{7}(r_{1}(e_{1}))$, this indicates that the inserted representation carries transferable semantic information corresponding to the bridge entity.
We apply this patching procedure across multiple layers and token positions, with three settings that differ in both the source of the intermediate entity and the model’s training stage: (1) Phase II with ID-derived intermediate entities, (2) Phase III with ID-derived intermediate entities, and (3) Phase III with OOD-derived intermediate entities. This alignment ensures that patching is conducted under conditions where the model is capable of reasoning over the relevant intermediate entity type. For completeness, we also report Phase I results, where patching yields negligible success, confirming that reasoning-relevant representations emerge only after generalization.
@@ 3.2 Explicitly Decodable $\neq$ Implicitly Informative
Having located the positions encoding intermediate entities, we next ask whether their internal role can be explained using existing interpretability tools. A key assumption in prior work is that if a hidden state encodes an intermediate entity, it should be decodable into a human-interpretable token, for example via the logit lens [lens2020, li2024understanding, yang2024large1, sakarvadia2024towards, zhang2024locate]. We test this assumption by measuring the decodability of $\mathbf{h}_{\texttt{r}_{1}}^{5}$, and examining whether decodability aligns with the emergence of reasoning behavior across training phases.
@@ 3.3 Geometric Regularity of Intermediate Representations via Cosine Lens
Most prior work focuses on decoding representations, but we take the reverse approach: given a known intermediate entity, can we identify recurring structure in how the model represents it across contexts to achieve consistent representations [wang2025reversal]? This reverse mapping is enabled by our earlier analysis in Section 3.1, which identifies the position $\mathbf{h}_{\texttt{r}_{1}}^{5}$ encoding the intermediate entity. At this anchor, we collect hidden states from queries sharing the same intermediate entity (Figure 1), and examine whether these vectors reflect a consistent internal pattern.
To assess consistency, we focus on structural alignment in the model’s embedding space, using cosine similarity—a common metric for semantic proximity in high-dimensional representations[goloviznina2024ve, rabinovich2023predicting, jurayj2022garden]. This allows us to examine whether the model reuses internal abstractions through representational geometry, instead of relying on explicit decodability.
Case Study: Visualizing Representational Clustering. To gain an initial sense of the representational patterns that emerge during training, we visualize the hidden states of a randomly selected intermediate entity that appears in multiple two-hop queries. For this entity, we extract $\mathbf{h}_{\texttt{r}_{1}}^{5}$ across relevant queries instances (where the intermediate entity is either ID-derived or OOD-derived) and compute pairwise cosine distances (defined as 1cosine similarity1-$\text${cosine similarity}) between them. We then project these high-dimensional vectors into two dimensions using Multidimensional Scaling (MDS)
@@
Tracking these scores, we find that the ID Cohesion Score rises steadily, aligning with Test-II generalization, while the OOD Alignment Score starts increasing later, following the rise in Test-OI performance (Figure 2). This suggests that successful implicit reasoning relies on representational consistency across diverse contexts: only when hidden states for the same entity align closely in cosine space can they be consistently reused for multi-hop inference. Thus, cosine-space clustering emerges as the model’s internal mechanism for semantic abstraction and generalization.
@@ 4 Closing the Loop: Explaining Behavioral Phenomena
Drawing on the mechanistic evidence in Section 3.3, we revisit our empirical observations in Section 2 and explain how they emerges from the internal dynamics of representation formation.
@@ 4.3 Why the First Hop, and Only the First Hop, Generalizes to OOD?
$\text{OOD-derived token}(e_{2})\xrightarrow{pulled\ by}\text{ID-derived token}(e_{2})\xrightarrow{align\ with}\text{latent}(e_{2}),$
@@ 5 Discussion
Our study, conducted in a controlled symbolic dataset environment, reveals key insights into the mechanisms of implicit reasoning in transformers, highlighting specific patterns and behaviors that clarify how multi-hop implicit reasoning emerges. These findings may provide valuable answers to existing questions about the implicit reasoning capabilities of LLMs. For instance, our observation regarding the requirement for query-level match offers a potential explanation for why knowledge learned from single-hop tasks does not easily transfer to multi-hop reasoning in LLMs [balesni2024two, zhang2024co, xu2025memorizing]. However, it is important to note that LLMs operate with far richer and more complex knowledge bases, and their internal knowledge interaction mechanisms likely differ from those in our controlled environment. Therefore, while our findings offer useful insights, they should be regarded as preliminary guidance rather than a complete explanation of the reasoning dynamics in LLMs.
Acknowledgement
This work is supported by Beijing Natural Science Foundation (L243006), National Natural Science Foundation of China (62476150), and the Tsinghua University (Department of Computer Science and Technology)-Siemens Ltd., China Joint Research Center for Industrial Intelligence and Internet of Things (JCIIOT).
@@ Appendix A Related Work
The initial focus on the mechanisms of implicit reasoning arose from the discovery that traditional single-hop knowledge editing methods are ineffective in the context of multi-hop implicit reasoning [zhong2023mquake, zhang2024enhancing, ju2024investigating]. This phenomenon since gained more significant attention, prompting researchers to investigate the mechanisms underlying implicit reasoning. Some works suggest that the failure of implicit reasoning is due to the intermediate entities not being properly processed [sakarvadia2024towards, yang2024large1]; Other works suggest that the reason lies in the intermediate entities being processed, but not being passed to the correct position [li2024understanding, yao2025cake, hopping-too-late]. The parallel exploration paths within the model are also a research direction [shalev2024distributional, wang2024understanding]. Recent works have found that the different knowledge composed as different hops used in implicit reasoning is stored in different layers of the model [zhang2024locate, yao2025cake]. Additionally, some studies propose that models rely on shortcuts to successfully complete implicit reasoning [yang2024large2].
@@
Some works have found that the conditions for models to learn implicit reasoning are quite demanding. Some have discovered through custom symbolic datasets that models generalize implicit reasoning abilities only after grokking, and that they require highly compositional data  [wang2024grokking]. Other works have found that simply providing optical knowledge is insufficient for enabling models to perform implicit reasoning, it requires training with corresponding multi-hop reasoning data [xu2025memorizing, zhang2024co, zhu2024towards]. Moreover, it has been observed that when the two pieces of foundational knowledge used for 2-hop implicit reasoning appear in different paragraphs of the training corpus, the model struggles to combine them for implicit reasoning. However, when these pieces appear within the same paragraph, the model’s accuracy in answering the corresponding 2-hop queries significantly improves [balesni2024two].
@@
As for the probing tools, most of the work uses decoding-based methods to probe intermediate entities. Many works assume that if a model encodes an intermediate entity, it should be able to explicitly decode it. They use the logit lens [lens2020]. as evidence of the presence of intermediate entities in implicit reasoning [sakarvadia2024towards, yang2024large1, li2024understanding]. A few studies have also trained a linear transformation layer to decode intermediate entities [shalev2024distributional].
@@
Although current Chain-of-Thought (CoT) based reasoning models have shown impressive performance, some perspectives argue that the thought chains do not truly reflect the model’s reasoning process [ma2025reasoning, anthropic_reasoning_models_2025]. Some works have begun to explore Chain of Thought (CoT) methods that do not verbalize intermediate steps. One approach is to replace the Chain of Thought (CoT) with dots, adjusting the number of dots based on the original length of the thought chain [pfau2024let, bharadwaj2024understanding]. Another approach to avoid verbalizing is to bypass decoding the tokens in the CoT process and directly use the last hidden state from the previous step as the input vector for the next step [hao2024training]. These approaches have achieved results comparable to traditional CoT on specific tasks.